### v5

# A Literature Review of Literature Reviews in Pattern Analysis and Machine Intelligence

Abstract
The rapid advancements in Pattern Analysis and Machine Intelligence (PAMI) have led to an overwhelming expansion of scientific knowledge, spawning numerous literature reviews aimed at collecting and synthesizing fragmented information. This paper presents a thorough analysis of these literature reviews within the PAMI field, and tries to address three core research questions: (1) What are the prevalent structural and statistical characteristics of PAMI literature reviews? (2) What strategies can researchers employ to efficiently navigate the growing corpus of reviews? (3) What are the advantages and limitations of AI-generated reviews compared to human-authored ones? To address the first research question, we begin with a narrative overview to highlight common preferences in composing PAMI reviews, followed by a statistical analysis to quantitatively uncover patterns in these preferences. Our findings reveal several key insights. First, fewer than 20% of PAMI reviews currently comply with PRISMA standards, although this proportion is gradually increasing. Second, there is a moderate positive correlation between the quality of references and the scholarly impact of reviews, emphasizing the importance of reference selection. To further assist researchers in efficiently managing the rapidly growing number of literature reviews, we introduce four novel, real-time, article-level bibliometric indicators that facilitate the screening of numerous reviews. Finally, our comparative analysis reveals that AI-generated reviews currently fall short of human-authored ones in accurately evaluating the academic significance of newly published articles and integrating rich visual elements, which limits their practical utility. Overall, this study provides a deeper understanding of PAMI literature reviews by uncovering key trends, evaluating current practices, and highlighting areas for future improvement.
Index Terms:
literature review, PAMI, umbrella review, AI-generated review, bibliometrics
1 Introduction
THE entropy of almost all natural and artificial systems in the universe exhibits a continuous increase. This principle extends to humanity’s knowledge system, which similarly experiences a steady rise in entropy. Such growing disorder within the knowledge system can lead to redundant efforts, hindering the efficient creation of new knowledge. Analogous to the role of gravity in shaping the early universe filled with diverse particles, the literature review serves as a critical mechanism for organizing dispersed knowledge within the academic sphere. Essentially, a literature review is a scholarly composition that not only demonstrates an understanding of the academic literature on a given topic but also situates this knowledge within a broader context. It compiles the most relevant and significant publications related to a specific research area, thereby offering a comprehensive overview of that field.
Nearly every field features its own body of literature reviews, particularly in the rapidly evolving domain of pattern analysis and machine intelligence. As a foundational technology and research direction for AI, it spans multiple areas such as image classification [21, 77, 73, 74], image segmentation [80, 85, 110, 39], object detection [142, 24, 98, 59], natural language processing [84, 40, 75], etc. As reported by the AI Index Report [78], there has been a striking surge in artificial intelligence publications, soaring from 200,000 in 2010 to nearly 500,000 by 2021. This exponential increase has subsequently led to a proliferation of related literature reviews. This trend is clearly illustrated in Fig. 1, which shows a marked increase in the annual publication of reviews, underscoring the growing prevalence of literature reviews in the field.
Figure 1: Annual publication trends of literature review in the field of PAMI. A notable rising trend can be observed since 2015, reflecting an increasing scholarly focus and growing recognition of the importance of review articles in synthesizing the state of research in PAMI. The data is collected from the Google Scholar search engine.
The increasing number of literature reviews across various AI subfields has raised several concerns. First, from the authors’ perspective, the characteristics of literature reviews in the PAMI field remain underexplored, including aspects such as structural conventions, statistical patterns, and adherence to established review methodologies. Second, researchers face difficulties navigating the rapidly expanding body of literature reviews, as the sheer volume within the same field makes it harder to identify which reviews are most relevant to their research objectives. Lastly, with the growing prevalence of AI-generated reviews, it becomes crucial to assess their benefits, limitations, and overall reliability compared to traditional human-authored reviews.
1.1 Research Question and Contribution
This paper specifically focuses on reviews within the field of PAMI and tries to answer the following three core Research Questions (RQs):
RQ1: What are the prevalent structural and statistical features shared among literature reviews within the PAMI field? Such information helps researchers develop a more specific understanding of literature reviews, thereby advancing the field’s overall comprehension of the role and significance of these reviews.
RQ2: What strategy can researchers employ to navigate the expanding corpus of literature reviews in PAMI more efficiently? A suitable strategy facilitates readers in navigating a large number of similar reviews and identifying the reviews that best match their interests and objectives.
RQ3: What are the advantages and limitations of AI-generated reviews, and what are the key differences between AI-generated and human-authored literature reviews in PAMI? A comprehensive understanding of AI-generated reviews enables users to benefit from the convenience offered by automated review technologies while minimizing the risk of potential misuse.
To address these questions:
  • We conduct both narrative review and statistical analysis to investigate the writing styles, meta-data distributions, and features of literature reviews in the PAMI field. This information aims to help researchers develop a more in-depth understanding of the nature of literature reviews.
  • We propose four real-time, article-level, and field-normalized bibliometric indicators for the quantitative evaluation of literature reviews. It is believed that the integrated use of these indicators may help researchers make more informed choices when selecting review articles.
  • We evaluate several existing automated review generation systems and conducted a systematic assessment and bibliometric analysis of these AI-generated reviews. The findings highlighted notable differences between AI-generated reviews and those written by humans, particularly in the quality of references, and the richness of visual elements.
In summary, we present a literature review of literature reviews and discuss the common concerns faced by existing reviews in the PAMI field. To the best of our knowledge, there has been limited scholarly attention to addressing these considerations. All the data and code framework used in this paper are publicly available at https://sway.cloud.microsoft/2TXEuPuNlDKEmC9p.
1.2 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 briefly introduces the criteria for literature screening, methods for the RiPAMI construction, and formulas for metric calculations. Section 3 provides a narrative analysis of the structure and content features of the selected reviews. Section 4 presents a statistical analysis of meta-data and content features of reviews in the RiPAMI. Section 5 further discusses how the proposed metrics can assist in efficiently selecting the proper reviews. The characteristics of human-authored versus AI-generated literature reviews are discussed in Sec. 6. In Sec. 7, we explore the challenges and future prospects of literature reviews. Finally, the paper concludes in Sec. 8.
2 Methodology
2.1 Scope
This study examines all fourteen types of review articles as classified in Grant’s comprehensive typology [33], including narrative review, systematic review, state-of-the-art review, etc. While this research is technically a tertiary or “umbrella” study, we opted for the more commonly understood term “literature review” to improve clarity and acceptance within the PAMI field. In addition, although the quantity of reviews is considerably smaller compared to that of normal papers, it remains impractical to analyze reviews within every field. Therefore, this paper will focus only on reviews in the PAMI field.
2.2 Literature Selection Criteria for Classical Review
To identify the features necessary for statistical analysis, we conduct a classical narrative review to determine the key characteristics of literature reviews in the PAMI field. This review utilizes resources such as Google Scholar, Semantic Scholar, IEEE Xplore, and ScienceDirect, covering publications from major publishers including IEEE, Elsevier, Springer, and others. To ensure that the selected papers are both review articles and highly relevant to the PAMI domain, we adopt a straightforward and effective approach. First, the 106 search keywords are derived from the scopes of prominent journals and conferences in the PAMI field, including terms such as speech recognition, optical character recognition, and self-supervised learning. Second, a filtering criterion is applied: the paper’s title must include either “survey” or “review”, and the specified search keyword must appear in the title or abstract. Finally, to maintain the quality of the included reviews, pre-prints not published in journals or conference proceedings are excluded, ensuring that all selected articles are peer-reviewed. This approach helps to maximize the inclusion of high-quality and relevant literature for subsequent analysis.
2.3 Literature Selection Criteria for Statistical Analysis
The immense number of literature reviews in the PAMI field emphasizes the need for approaches beyond manual selection, analysis, and synthesis. Therefore, to ensure a comprehensive and accurate analysis of literature reviews in the PAMI field, we develop an automated process that simulates the manual selection procedure to construct the RiPAMI database.
2.3.1 Data Source
TABLE I: Comparisons between various sources: “Counts only” in the citations column means that the source only records the citation counts, while “Complete” signifies that the data source provides a complete list of citations.
Database Title & Authors Venue Abstract Citations References Source Types Free
arXiv $\times$ $\times$ $\times$ API-based
CrossRef Counts Only API-based
Google Scholar Complete $\times$ Crawler-based
IEEE Xplore Counts Only $\times$ API-based
Semantic Scholar Complete API-based
Web of Science Complete API-based $\times$
Scopus Complete API-based $\times$
arXiv Data File $\times$ $\times$ $\times$ Snapshots
CrossRef Data File Counts Only Snapshots
RiPAMI(Ours) Complete Snapshots
Reliable data sources for analyzing extensive reviews are fundamentally important. Based on the means of data acquisition and storage, existing scientific scholar data sources may be classified into two main categories: web-based and snapshot-based sources. Web-based source data refers to the meta-data that can be retrieved from the data provider in real-time with the use of the web crawler or the API (e.g. Semantic Scholar, arXiv, CrossRef). Such an approach would only consume a small amount of storage on the local machine but needs to query meta-data every single time. On the contrary, the offline snapshot consumes a larger amount of storage space but eliminates the need for frequent API queries. Since the snapshot is a mirror image of relevant papers before a specific time, its data remains unchanged over time compared to the API-based sources. As a result, the snapshot ensures a consistent dataset in all of the experiments, avoiding issues of irreproducibility caused by changes in the provided web retrieving service.
Table I compares various most commonly used data sources. However, none of these sources is perfect. For example, arXiv is a free distribution service and an open-access archive for millions of scholarly articles in various fields. However, the arXiv data source fails to provide the publication venue, citations, and references. Semantic Scholar seems promising, but it suffers a lower update rate than arXiv and a narrower search scope than Google Scholar. Google Scholar is an online search engine that indexes scholarly literature from a wide range of disciplines. Although Google Scholar uses powerful automated programs to retrieve files for inclusion in its search results, it still faces challenges in many areas. Beel [9, 10, 11] argues that Google Scholar places a high weight on citation counts in its ranking algorithm and has therefore been criticized for exacerbating the Matthew effect [83]. Moreover, the citation counts displayed on Google Scholar are subject to manipulation by complete nonsense articles indexed on Google Scholar (e.g. citations from AI-generated pre-print papers published on arXiv should have been ignored). Therefore, a promising engineering solution is to leverage the strengths of various approaches to overcome the weaknesses of each approach, as will be detailed in the next subsection.
2.3.2 Database construction
[…]
Figure 2: A schematic overview of the database construction process. From initial keyword selection to the final SQL-based RiPAMI snapshot, three key steps are implemented to ensure the data in RiPAMI is clean, accurate, and reliable. For simplicity, the polar diagram shows only a subset of the keywords used for retrieval. Since citation counts fluctuate over time, the retrieval date for related information is set to October 2024.
As illustrated in Fig. 2, we implement a structured process to construct the RiPAMI database, ensuring that the selected articles are review papers relevant to the pattern recognition field. The process begins with searching with keywords derived from the scopes of relevant journals and conferences. These keywords are then formatted into query strings for the arXiv API as follows: (ti:“review” OR ti:“survey”) AND (ti:“keyword.lower()” OR abs:“keyword.lower()”). This structure specifically retrieves articles whose titles or abstracts include both the terms “review” or “survey” and the specified keywords. Additionally, we apply regular expressions to check that the abstracts of returned articles include the relevant keywords, ensuring that only appropriate review articles enter the RiPAMI database. To avoid potential copyright and licensing issues, papers are retrieved and downloaded using the arXiv’s API. Calls to the API are made by means of HTTP requests to a certain URL. The responses will be parsed and stored in an SQL-based database. Through this process, we collect a total of 4,106 samples. To further refine our database, we employ a double-checking phase. This involves a GPT-based filter to preliminarily classify papers, followed by a manual validation step to remove any noisy or irrelevant entries. Lastly, we supplement the remaining 3099 entries with additional metadata, including the publication date, citation counts, and reference details. As mentioned in Sec. 2.3.1, we suggest enriching the meta-data of papers by leveraging a combination of disparate data sources. Considering the potential legal risks of crawling to obtain academic data from Google Scholar, Semantic Scholar API was employed to obtain additional meta-data, such as citation and reference details which are not provided by the arXiv API. The final SQL-based snapshot encompasses a wide range of data, including ID, title, publication date, citations, and more, facilitating efficient data retrieval and statistical analysis.
2.4 Bibliometric Indicators for Reviews
Literature reviews reflect the state of their respective fields, highlighting current research priorities and offering a glimpse into emerging trends. Despite their significance, systematic methods for evaluating literature reviews remain underdeveloped. By introducing quantitative tools, bibliometric methods offer a promising solution, enabling researchers to objectively analyze literature reviews and uncover key patterns and trends. Bibliometrics is a research field focused on the use of quantitative methods and statistical analysis to evaluate various aspects of scholarly publications [15]. These methods and analysis, including widely recognized metrics like citation counts and impact factors, offer researchers valuable insights in their daily scientific work, aiding them in making prompt and informed decisions. However, most existing bibliometric methods suffer from several limitations including unfair comparison, misuse, and manipulation as introduced in the “San Francisco Declaration on Research Assessment” (DORA) [30] and the Leiden Manifesto [41]. Notably, many of these metrics were not specifically designed for evaluating literature reviews, further limiting their applicability in this context.
In light of these limitations, we propose two new impact indicators, $TNCSI$ and $IEI$, along with two quality indicators, $RQM$ and $RUI$.
2.4.1 Topic Normalized Citation Success Index (TNCSI)
Most existing bibliometric methods face challenges in conducting cross-disciplinary comparisons due to variations between fields [117, 97, 109, 115]. Although a few methods offer the capability for cross-disciplinary comparison, they either require manual assignment of field-specific keywords or involve complex and costly retrieval to calculate metrics. These drawbacks have, to some extent, hindered individual researchers from adopting these methods, thereby limiting the broader application of the metrics.
To fairly compare the external influence of literature reviews across different fields within an acceptable cost, we propose guiding Large Language Models (LLMs) to generate the topic key phrase and then calculating the success index of the current review within the same topic (see Appendix A.1 for details). The $TNCSI$ estimates the impact of research publications within an LLM-generated topic by normalizing citation counts to a scale between 0 and 1. The formula for calculating the $TNCSI$ is:
[…]
where $\lambda$ is the scale parameter obtained through maximum likelihood estimation. Detailed steps for estimating $\lambda$ and constructing the probability density function are provided in the appendix.
The $TNCSI$ demonstrates favorable mathematical properties and interpretability. First, the $TNCSI$ algorithm employs maximum likelihood estimation to convert the probability mass function into a probability density function. This process ensures that, in theory, the $TNCSI$ differentiates between papers with distinct citation counts, avoiding the assignment of identical values to them. Second, the $TNCSI$ possesses physical significance, representing the probability that a specific paper’s citation count surpasses that of any other paper on the same topic. For example, a paper with a $TNCSI$ of 0.5 means it has more citations than half of the papers within the same topic. Furthermore, the calculation of $TNCSI$ is computationally efficient, as it avoids requiring a complete or precise ranking of all related papers. Instead, it provides a reliable estimation of the probability that the given paper surpasses others in terms of citation count within the same topic.
2.4.2 Impact Evolution Index (IEI)
Imagine a scenario where two papers, A and B, receive the same number of citations. The number of new citations per month for A remains steady, whereas the number of new citations for B grows exponentially. In this context, while acknowledging the importance of Paper A, it is generally assumed that Paper B holds a greater reference value. Analyzing the popularity or citation trends of the literature may help researchers stay informed about the latest developments and identify potential areas for future research.
The $IEI$ is defined as the average slope across $l$ distinct points on a $n=l-1$ degree Bézier curve representing the citation trend over time. The formula for calculating $IEI_{L_{l}}$ is:
$IEI_{L_{l}}=\sum_{a=0}^{l-1}\frac{(y_{i}/x_{i})}{l},$ (2)
where $l$ represents the number of months observed. $x_{i}$ and $y_{i}$ are the components of the tangent vector at the $a$-th point on the Bézier curve, indicating the magnitude along the $x$- and $y$-axes, respectively. More details can be found in Appendix A.2
With the guidance of the $IEI$, researchers may further discern the various literature reviews within the same field that exhibit close citation counts. A higher $IEI$ indicates that an increasing number of studies reference the review, signaling its growing influence and attention. Therefore, in practical applications, researchers may prioritize the study of more promising reviews, as indicated by higher positive $IEI$ values, to mitigate redundancy and improve efficiency.
2.4.3 Reference Quality Measurement (RQM)
A literature review, in its essence, can not fabricate insights from a void. It fundamentally relies on the substance of existing references. Without a solid foundation of credible and high-quality sources, a literature review may lack the necessary building blocks to construct a meaningful analysis or argument. These sources provide the empirical evidence and theoretical context that ground the review, making the role of references indispensable in the creation of a substantial literature review.
The RQM incorporates both the quality and timeliness of references by modifying the Gompertz function, which exhibits a sigmoidal growth pattern—slow at the beginning and end, with a rapid increase in the middle phase:
$RQM=1-e^{-\beta\cdot e^{-(1-ARQ)\cdot S_{mp}}}.$ (3)
In this equation, $\beta$ represents the shift parameter, $ARQ$ denotes the average reference quality, while $S_{mp}$ refers to the median semester count of the reference age, indicating the time from the publication date of the references to the issuance of the review (which will be detailed in the Appendix A.3).
Beyond simply assessing the average impact of references, the $RQM$ integrates references’ timeliness and its influence on review quality. This metric addresses the tendency among authors to rely heavily on classic works, which may no longer fully align with current research trends. Additionally, the adjustment coefficient in $RQM$ is calibrated through statistical and optimization algorithms, tailored to the unique characteristics of each research discipline. This data-driven approach minimizes reliance on heuristic parameter settings and increases RQM’s adaptability across various academic disciplines.
2.4.4 Review Update Index (RUI)
The $RUI$ refers to the measure of the extent to which a literature review is required to be updated due to the iteration of technology, theory, etc. The index is related to both the literature itself and the research interests of the topic. Generally, a high update index suggests that a literature review is in need of an immediate update. Conversely, a lower update index implies that few advances have been made to the investigated field and the review is still up-to-date.
The $RUI$ quantifies the necessity of updating a literature review by combining two indicators: the Coverage Difference Ratio (CDR) and the Review Aging Degree (RAD). The formula for calculating the $RUI$ is:
$RUI=p\cdot CDR+q\cdot RAD.$ (4)
Here, $p$ and $q$ are weighting coefficients, set to 10 and 5, respectively. Detailed definitions and calculations of the $CDR$ and $RAD$ are provided in Appendix A.4.
The proposed $RUI$ integrates both the popularity of the research field and the natural aging of individual reviews. It reflects how reviews require updating over time while taking into account the pace of progress within the field. By balancing these factors, $RUI$ offers real-time insights into how urgently a review needs updating. This urgency may arise from rapid developments in the field or the natural aging of the review’s content. $RUI$ moves beyond the simplistic reliance on publication dates, providing a more contextualized assessment of its timeliness.
2.5 Information Extraction for Reviews
Given that this study aims to identify the common features of reviews within the field, solely relying on manual inspection would be both impractical and resource-intensive. To further investigate the features of reviews in the PAMI field, we employed LLM-based information extraction techniques to analyze the review content and parse the response to structured results. The proposed approach converts unstructured text into structured data, facilitating further analysis and interpretation. Details regarding the information extraction method can be found in Appendix B.
3 Classical Review of Reviews in PAMI
This chapter provides a classic, narrative overview of the PAMI field, aiming to identify common features shared by existing reviews and laying the foundation for the subsequent statistical analysis.
3.1 Popular Structure of Reviews in PAMI
The review’s structure is alternatively referred to as the framework of the review. It is the outline that authors consider when starting to conduct the survey. A well-designed structure is believed to enhance the paper’s readability and facilitate the reader’s comprehension of the paper’s concepts and knowledge. Typically, this framework encompasses foundational parts or sections such as the abstract, introduction, methodology, discussion, and conclusion, each fulfilling a specific function.
3.1.1 Abstract
The abstract is the most important part of a paper, as it typically summarizes the core content, helping readers quickly grasp the main points and conclusions of the article. An abstract may vary in its level of detail, but a structured abstract improves clarity by dividing the content into clearly labeled sections, enabling readers to quickly understand the main aspects of the study. A structured abstract consists of labeled sections such as Background, Objective, Methods, Results, Conclusion/Discussion, and possibly Limitations [55, 12]. For example, Sariyanidi et. at. [103] structured their abstract by providing the background on facial affect, outlining the persistent challenges in the field, and summarizing their research methods, results, and discussion.
3.1.2 The Beginning Part
Similar to the research article, the introduction section of the literature review usually includes contextualizing the research topic, identifying the knowledge gap, defining the scope and objectives, and laying the foundation knowledge for wider audiences. “Introduction” usually lies at the very beginning of the paper. Most reviews first introduce the definition of the topic to acquaint readers with a basic understanding of the field. For instance, the term “Named Entity Recognition” may be unfamiliar to many scholars outside the field of natural language processing. A clever introduction might provide the origins of the terminology and inform the readers what is named entity recognition, as has been done in Ref [88]. A brief definition of the field in the introduction is also welcomed for some relatively popular fields, e.g., the literature review [102] examines the concept of image categorization in the first sentence.
In addition to contextualizing the research topic, many introductions also highlight the existing research and identify gaps or challenges in the current understanding of the topic. It sets the stage for the literature review by explaining why the research being reviewed is necessary and what gaps are expected to be filled. Wang et al. [123] emphasize that while there are comprehensive surveys of self-supervised learning in computer vision, there is a lack of a similar overview specifically tailored to the remote sensing community.
Readers need to evaluate if the article is worth reading before investing more time in it. A statement of the scope or contributions of the article is a way for readers to quickly assess its relevance. In the first section, Wangkhade et al. [125] provide us with several important contributions including analyzing well-known technologies, proposing taxonomies of approaches, and summarizing benefits and challenges of sentiment analysis.
Preliminaries and problem formulations are also popular with readers, as they both provide profound background knowledge. Given that the Gumbel-max involves considerable mathematical concepts and calculations, a “Preliminaries” section in the review of the Gumbel-max trick [44] serves the purpose of providing background information and basic understanding related to the Gumbel-max trick.
For systematic reviews, a clear statement regarding the review methodology (e.g., literature search strategy and criteria for inclusion and exclusion) is essential [92, 54]. Such a statement helps to minimize potential bias and enables other researchers to replicate the study’s findings. Memon et. at. [82] dedicate a section to detailing their review process, including subsections on the review protocol, inclusion and exclusion criteria, search strategy, study selection process, quality assessment criteria, and data extraction and synthesis, ensuring transparency and rigor in their methodology.
3.1.3 The Middle Part
The middle part of the survey paper, also known as the review part, presents a detailed examination of the relevant research studies, methodologies, findings, and theories related to the chosen theme of the review. Beyond that, The middle part synthesizes the information from related studies to provide a cohesive and integrated understanding of the topic. This synthesis not only aligns the disparate studies but also evaluates their contributions and the interrelations among them. It often includes comparative analysis, highlighting similarities and differences in approaches, and may identify limitations in the existing research that warrant further investigation.
To understand more comprehensively how the middle part is organized and synthesized, we examine a large number of review papers in the field of PAMI and propose a typology for organizing the review section based on these analysis. By analyzing the writing styles and structures of the review part, we categorize them into three main types: method clustering, challenge-oriented, and others.
Similar to the mapping or scoping review, the method clustering-based review part refers to grouping methods according to their technical characteristics and presenting them in separate sections or subsections. Authors need to identify different sections depending on the topic of the article and arrange the same type of methods as closely as possible. Featuring a clear and well-organized article structure, reviews with method clustering-based parts are widely favored by researchers. The reader can gain a comprehensive understanding of the typology, details, and advancements of technology within a specific field. However, such a type of literature review is less tailored for readers who are completely unfamiliar with the subject. This is because readers who lack basic knowledge of the field may not understand the connection between the methods mentioned in different sections, or are even confused to select a proper algorithm for a certain task. By way of illustration, Minaee et al. conducted a comprehensive survey on deep learning-based image segmentation models [85] in a method clustering-based manner. They grouped the models into 10 categories based on the adopted model architectures, such as fully convolutional models, encoder-decoder-based models, and attention-based models. In this way, the reader can quickly gain an understanding of the typologies of deep learning-based image segmentation models and how they differ from each other.
Known for its practicality, the challenge-oriented review part focuses on presenting research methods and findings that address a specific challenge or task. For this type, each section or subsection focuses on a specific challenge or task and centers the discussion on how to address that challenge. The references cited within each section are primarily directly related to addressing that challenge to provide support and evidence. Challenge-oriented reviews are well-suited to application-oriented fields, where readers at different levels of expertise can easily and quickly seek out appropriate solutions within the literature to address the problem they are facing. Despite the advantages, challenge-oriented reviews may fail to provide a comprehensive elaboration of fundamentals and methods details. Bandini et al. divided their survey about egocentric vision hand analysis into several sections including hand segmentation, hand detection, and hand identification [8]. Each section contains multiple subsections, and each of the subsections investigates a certain challenge. (e.g. subsections entitled “Robustness to Illumination Changes” and “Lack of Pixel-Level Annotations” were arranged in the “Hand Segmentation” section ).
In addition to those two types of reviews, a hybrid middle part strikes a fair balance between comprehensiveness and practicality. Typically, the hybrid middle part refers to a challenge-oriented framework incorporating a method clustering-based sub-framework. One advantage of a hybrid review is that it simultaneously offers multiple solutions to a certain problem along with the classification of the respective methodologies. An example is the study carried out by Guo et al. [37] in which authors divide sections by tasks and introduce similar approaches group by group.
It is worth noting that the primary goal of the typology is not to exhaustively categorize every type of review but to help readers and automated review generation systems better understand the common review structures within the PAMI field. Consequently, certain specialized types of reviews, such as meta-analysis, may fall outside the scope of our typology.
3.1.4 The Ending Part
The subsequent ending part serves as a succinct conclusion to the reviewed studies. Within this section, the authors prefer to sum up the key findings to answer the question in the beginning part. It usually highlights both the advantages and disadvantages of reviewed literature to conclude research gaps and suggest potential avenues for future research.
One of the main objectives of the concluding part is to provide a relatively concise summary of the main insights derived from the reviewed studies and highlight their significance and relevance to the research topic. Some literature reviews conduct comprehensive analysis and synthesis, which in turn run up to a considerable number of pages. The paper [63] consists of 51 double-column pages in the main body, which poses a challenge for readers who may not have the time to peruse it thoroughly. Fortunately, this paper also includes a concise summary which provides readers with an efficient understanding of the content and allows them to navigate to relevant sections of interest.
When discussing further in the ending part, the authors may attempt to identify gaps and point out future directions through the analysis of existing literature. Some of the papers present the gaps and future directions separately. Hassain et al. [45] first discussed major challenges of multi-view video summarization such as lack of synchronization, instability of camera, and crowded scenes individually. Where there are challenges, there are future research directions. In addition to challenges, the authors also provide recommendations and future directions from various perspectives including models, benchmark datasets, and agents-based MVS, etc. Conversely, some papers choose to combine discussions of current issues with emerging research trends, as seen in the ’Future Directions’ section of Ref[135].
For literature reviews delving into pragmatic methodologies, the inclusion of an applications section is quite fitting. Ref [60] offers a comprehensive review of various surveys on convolutional neural networks, dedicating a distinct section to the typical applications of 1-D, 2-D, and multidimensional CNNs. Similarly, Ref [71] illustrates the deployment of few-shot learning techniques across disciplines like computer vision, natural language processing, and reinforcement learning.
3.2 Review of Reviews in Various Sub-fields in PAMI
In this section, we offer a subjective analysis of reviews from diverse fields. Through this narrative approach, we aim to examine the composing characteristics shared by reviews across various fields.
3.2.1 Computer Vision
Computer vision is one of the most popular sub-fields of pattern recognition and machine intelligence. This section will focus on the literature reviews within the realm of computer vision.
Image Classification refers to the task of assigning a label or a category to an input image, which is one of the most renowned tasks in the field of computer vision. A survey by Rawat [102] explores the development and advancements of deep convolutional neural networks (CNNs) in the field of image classification. The paper covers the historical context, their role in the deep learning renaissance, and the notable contributions and challenges faced in recent years. It highlights the remarkable progress of CNNs in image classification, while also acknowledging the ongoing research efforts to address challenges and provide recommendations for future exploration. Schmarje et al. [104] provides a comprehensive survey on semi-, self-, and unsupervised learning methods for image classification. The survey compares and analyzes 34 different methods based on their performance and commonly used ideas, highlighting the trends and research opportunities in the field. Through comprehensive analysis, the authors reveal the potential of semi-supervised methods for real-world applications and identify challenges such as class imbalance and noisy labels. Furthermore, the paper emphasizes the importance of combining different techniques from various training strategies to improve overall performance. In addition to CNNs, there exist alternative techniques for image classification. The paper [20] presents a comprehensive analysis of Support Vector Machines (SVM) in image classification. It discusses various techniques that can enhance classification accuracy and highlights its advancements. Liu et al. [69] investigate more than 100 different visual Transformers comprehensively in three fundamental CV tasks including classification, detection, and segmentation. They also propose a taxonomy to categorize various transformers into six groups.
Object Detection entails identifying and localizing objects of interest within an image or video. Liu et al. [63] offers a comprehensive survey on the advancements in deep learning-based generic object detection. This paper discusses an extensive range of issues, including detection frameworks, taxonomies, feature depiction, training strategies, and evaluation metrics. Though there have been significant advancements in generic object detection, the detection of small objects, which focuses on identifying objects with a small size, still presents challenges. The review conducted by Cheng et al. [24] investigates 181 papers, constructs two large-scale datasets (SODA-D and SODA-A), and evaluates the performance of mainstream small object detection methods. Object detection demonstrates the utility and effectiveness across multiple domains. Li et al. [59] and Litjens et al. [62] investigate numerous methods and applications of object detection in remote sensing and medical image analysis respectively, showing that these methods have the flexibility to be applied in various scenarios and meet different needs.
Image Segmentation is the process of dividing an image into meaningful and distinct regions to facilitate analysis and understanding. As described earlier, Minaee et al. [85] proposed a taxonomy for image segmentation methods which divides models into 11 categories. In addition to the taxonomy, the authors evaluate the quantitative performances of various methods on popular benchmarks. The paper also identifies open challenges and proposes promising research directions for future advancements in deep-learning-based image segmentation. Given that most image segmentation algorithms heavily rely on expensive pixel-level annotations, interest in weakly supervised image segmentation methods has increased. Ref [108] surveys label-efficient deep image segmentation methods. According to the paper, weakly supervised segmentation approaches can be categorized into four hierarchical types, ranging from no supervision to inaccurate supervision. The authors investigated each of these four methods in separate sections, highlighting the strategies used to bridge the gap between weak supervision and dense prediction. Image segmentation techniques have a wide range of applications in the field of medical image processing, as introduced in [100, 133, 110, 72].
3.2.2 Natural Language Processing
Acclaimed as the jewel of the artificial intelligence crown, natural language processing (NLP) stands as a pivotal domain within the field of PAMI. Here, we provide a further discussion on several popular NLP research directions.
Named Entity Recognition (NER) involves identifying and classifying named entities in text, such as person names, organizations, locations, and dates. The survey by Li et al. [58] begins by introducing NER resources, including tagged NER corpora and off-the-shelf NER tools. Then, authors categorize existing works based on a taxonomy that considers distributed representations for input, context encoder, and tag decoder. The paper surveys representative methods for applying deep learning in various NER tasks, and provides a valuable reference for designing deep learning-based NER models. NER serves as the foundation technique for various natural language applications, such as relation extraction [90], knowledge graph [1], etc. Due to the linguistic variance of different languages, NER methods may also vary from language to language. Surveys about various language-specific NER could be found in [64, 126, 99].
Sentiment Analysis focuses on determining the sentiment or emotion expressed in text, such as positive, negative, or neutral. Yadav et al. introduce the process of gathering and analyzing people’s opinions and sentiments from various sources such as social media platforms and blogs in their paper [134]. The paper evaluates and compares different approaches used in sentiment analysis, with a focus on supervised machine learning methods like Naive Bayes and SVM algorithms. The common application areas of sentiment analysis and the challenges involved in accurately interpreting sentiments are also reported. The paper by Yue [140] categorizes and compares a large number of techniques and methods from three different perspectives: task-oriented, granularity-oriented, and methodology-oriented. It also explores different types of data and advanced tools for research, highlighting their strengths and limitations.
Language Modeling involves training models to understand and generate human language. As early attempts, recurrent neural networks achieved desirable performance and wide application at that time, despite some shortcomings. The paper [139] specifically focuses on RNNs and long short-term memory (LSTM) cells. The authors highlight the limitations of traditional RNNs and emphasize the significance of LSTM in handling long-term dependencies. They discuss various LSTM cell variants and their performance on different characteristics and tasks. Furthermore, the paper also categorizes LSTM networks into two major types: LSTM-dominated networks which optimize connections between inner LSTM cells, and integrated LSTM networks which incorporate advantageous features from various components. Recently, LLMs have drawn widespread attention. LLMs demonstrate significant performance improvements and unique abilities such as in-context learning, setting them apart from smaller-scale models. By investigating more than 600 works, Zhao et al. [149] conduct a comprehensive review of the recent advancements in LLMs. The authors discuss the evolution of language modeling techniques, from statistical models to neural models, and highlight the emergence of pre-trained language models as a powerful approach in NLP tasks. The survey focuses on LLMs with a parameter scale exceeding 10 billion and explores four key aspects: pre-training, adaptation tuning, utilization, and capacity evaluation. The paper also presents available resources for developing LLMs and discusses important implementation guidelines. Overall, this survey serves as an up-to-date and valuable reference for researchers and engineers interested in the field of LLMs.
3.2.3 Others
Reviews in other popular sub-fields will be investigated in this section.
The paper by Zhou et al. [151] covers the evolution of pre-trained foundation models from BERT to ChatGPT and highlights their significance as parameter initializations for downstream tasks. The survey explores popular pre-trained foundation models in text, image, and graph modalities, discussing their components, pre-training methods, and advancements thoroughly. The paper also addresses topics including model efficiency, compression, security, and privacy, while offering valuable insights into scalability, logical reasoning ability, and cross-domain learning. Another survey paper [138] presents a comprehensive review of the state-of-the-art in self-supervised recommendation (SSR). The paper proposes an exclusive definition of SSR and develops a taxonomy that categorizes existing SSR methods into four categories: contrastive, generative, predictive, and hybrid. It further introduces an open-source library called SELFRec, which incorporates a wide range of SSR models and benchmark datasets. Through rigorous experiments and empirical comparison, the paper derives significant findings related to the selection of self-supervised signals for enhancing recommendation. The conclusion highlights the limitations and outlines future research directions in the field of self-supervised recommendation.
3.3 Summary
From the discussion above, it is clear that in the field of PAMI, most authors exhibit similar preferences when composing reviews. Common practices include establishing taxonomies, highlighting future challenges and research directions, offering foundational background knowledge, summarizing the performance of various methods on standardized datasets, and examining the real-world applications of algorithms. These preferences will serve as the content features of the review and will be further statistically analyzed in the next section.
4 Statistical Analysis of Reviews in PAMI
In this section, we perform a statistical analysis of reviews in the RiPAMI database, with the goal of uncovering statistical patterns of reviews across various fields of PAMI.
4.1 Meta-data Statistics of Reviews in PAMI
The database consists of more than 3000 literature reviews from a variety of sources, publication years, and fields. To elucidate the characteristics of the RiPAMI database, we conduct a statistical analysis and plot the result in Fig. 3.
Figure 3: Statistics of the RiPAMI database. The samples in the RiPAMI database are characterized by a diversity of publication dates, scholar impact, reference numbers, etc., offering a comprehensive reflection of the state of reviews in the PAMI field.
Years of Publication Figure 3 (a) presents the distribution of publication years for literature reviews in the RiPAMI database. A clear upward trend in the number of reviews over time is evident, mirroring the pattern observed in Fig. 1. Both figures show a steady increase, with a notable surge between 2019 and 2020. This trend reflects a strong alignment between our sample data and the original dataset in terms of statistical consistency.
Number of the References The number of references in a survey paper may influence its credibility and reliability. As shown in 3 (b), the distribution of literature review references in the RiPAMI database follows a log-normal pattern.
Number of the Authors A review paper typically aims to provide a comprehensive overview of a specific topic. Involving multiple authors with diverse expertise could enhance the depth and breadth of the review. Fig. 3 (c) indicates that the majority of reviews are written by fewer than 10 authors.
Number of the Citations We count the citations of papers in RiPAMI and plot them in Fig. 3 (d). A power law distribution of received citations could be found. This phenomenon where a small subset of papers receives the majority of citations is sometimes referred to as the “Matthew Effect” or the “Pareto principle”, as reported in [83, 18].
Length of Reviews The words count and pages number of a review article can partially reflect its depth and breadth. In general, more detailed and comprehensive reviews tend to contain a higher word count and longer pages, thus requiring more time and effort to conduct and comprehend. As shown in Fig. 3 (e) and (f), review articles in the PAMI field typically range from 5,000 to 20,000 words and span 10 to 40 pages. At an average reading speed of 240 words per minute [17], reading a full review would take over 20 minutes.
Number of the Visual Elements Visual elements refer to visual representations of information, such as images and charts. They provide a clear presentation of data, simplify complex content, and enhance reader comprehension. By employing LLMs for information extraction, we may ascertain the number of figures and tables each review article contains, as depicted in Fig. 3 (g) and (h). It can be observed from our analysis that the majority of review articles contain fewer than 10 figures or tables.
Further statistical details, such as maximum (Max), minimum (Min), mean, median, and mode, are available in Tab. II.
TABLE II: Statistical Meta-data Summary of RiPAMI
Attribute Max Min Mean Median Mode
Pub Year 2024 2010 2021 2022 2024
Authors 59 1 4.61 4 3
Refs 1,700 1 138.19 120 66
Cites 11,578 0 78.48 11 0
Words 57,883 393 13,613.57 12,974 5,903
Pages 118 2 25.24 22 20
Figs 77 0 7.42 6 3
Tabs 47 0 4.38 3 1
4.2 Review Feature Statistics
As discussed in Sec. 3, review articles in the PAMI field exhibit shared features. In this subsection, we apply information extraction techniques to analyze the review articles within the RiPAMI database, aiming to explore the overarching statistical patterns and trends within PAMI reviews. Specifically, we introduce an ensemble query approach based on LLMs to extract key information from the raw content, determining whether a review (1) Taxonomy: explicitly outlines a taxonomy for existing methods; (2) PRISMA: includes a dedicated section specifying inclusion/exclusion criteria, similar to PRISMA guidelines [92]; (3) Preliminary: provides background knowledge in a distinct section; (4) Benchmark: presents tables or figures comparing quantitative performance across methods; (5) Application: discusses current applications within the field; (6) Discussion: includes a section addressing challenges, future directions, or limitations of the field. These six features are recorded in binary format (0/1) and are used for subsequent statistical analysis. Details regarding the extraction of these content features are provided in Appendix B.
Figure 4 presents our analysis of content features across a sample of over 3,000 review articles. We found that the majority of reviews tend to include a dedicated section discussing challenges, future directions, or opportunities within the investigated field. The high prevalence of discussion of future and existing challenges reflects established academic norms, where reviews are expected to identify unresolved issues and propose future research paths. Notably, the proportion of reviews adhering to PRISMA in the PAMI field remains below 20%, regardless of whether they have been peer-reviewed or not. This indicates that there remains considerable scope for the advancement of rigorous systematic reviews within the PAMI field.
Figure 4: Statistical Analysis of Review Features in the RiPAMI Database Using LLM-based Information Extraction Techniques: “Published” refers to peer-reviewed review papers, while “Non-published” denotes preprints without peer review.
Figure 5: Trends in the Proportions of Review Features over Time: The proportion of most features has been rising progressively over the years. For improved visual clarity, Gaussian smoothing is applied to the raw data.
Figure 5 illustrates the changing proportions of six key review content features over the years. Most of the analyzed features show a steady upward trend over time, with “Discussion” and “Preliminary” seeing the most notable increases. The proportion of “PRISMA”, which reflects the methodological rigor of reviews, has risen by over 10% between 2015 and 2024, indicating that reviews in the PAMI field are becoming more structured and methodologically sound.
4.3 The Importance of References
As the saying goes, ’You can’t make something out of nothing.’ The same principle applies to literature reviews: a well-crafted review is typically built upon a thorough analysis of high-quality references. To further explore the relationship between the quality of references in literature reviews and the review’s impact, we conduct a visual analysis of the distributions of $ARQ$ (representing the quality of references), $S_{MP}$ (representing the recency of references), and the corresponding $TNCSI$ (representing the scholar impact) for reviews published after 2020, as shown in Fig. 6.
Figure 6: Visualization of references quality of human-authored and AI-generated literature reviews. Each point represents a review sample. The $TNCSI$ exhibits a positive correlation with $ARQ$ and $RQM$, and a negative correlation with $S_{MP}$.
Figure 6 (a) highlights the correlation between $RQM$ and $TNCSI$. The scatter plot indicates a notable trend where lower $RQM$ values are generally associated with lower $TNCSI$ values, with a significant concentration of points observed in the range of $TNCSI$ between 0.0 and 0.2. Figure 6 (b) demonstrates that reviews citing higher-quality references are more likely to achieve higher academic impact. Generally, it is difficult for reviews with an $ARQ$ below 0.5 to reach a $TNCSI$ higher than 0.5. Furthermore, Fig. 6 (c) shows that reviews with a median reference age exceeding 10 semesters (i.e., 5 years) tend to exhibit relatively limited impact.
TABLE III: Correlation between Reference Quality and TNCSI. “Sig.” indicates significance, and ✓indicates the corresponding p-value less than 0.05.
Metric Pearson Pearson p-value Spearman Spearman p-value
$ARQ$ 0.43 $6.99\times 10^{-40}$ 0.42 $1.11\times 10^{-37}$
$S_{mp}$0.27 $6.29\times 10^{-16}$0.32 $2.21\times 10^{-21}$
$RQM$ 0.53 $1.34\times 10^{-62}$ 0.51 $1.86\times 10^{-57}$
To further quantify the relationship between reference features and cumulative academic impact, we present the Pearson and Spearman correlation coefficients between various indicators and $TNCSI$ in Tab. III. Both $ARQ$ and $RQM$ show moderate positive correlations with $TNCSI$, with Pearson and Spearman coefficients around 0.43 and 0.53, respectively. These correlations are highly significant, as indicated by p-values close to zero. In contrast, $S_{mp}$ exhibits a negative correlation, as indicated by Pearson and Spearman coefficients of -0.27 and -0.32, respectively. These statistically validated results demonstrate the importance of reference quality in influencing academic impact of reviews.
4.4 Insights into Field Trends via IEI and RUI
As outlined in Sec. 2.4.2 and 2.4.4, $IEI$ and $RUI$ are indicators used to evaluate changes in citation trends and the recency of individual articles. By utilizing these two indicators, we generate scatter plots depicting the temporal distribution of $IEI$ and $RUI$ values for review papers within the domains of recommendation systems and language modeling (as illustrated in Fig. 7).
Figure 7: Temporal Distribution of $IEI$ and $RUI$ in Surveys for Recommender Systems and Language Modeling: Higher $IEI$ and $RUI$ values in language modeling over the past two years indicate a trend of intensified development. Although surveys in popular fields tend to attract more academic attention, they also require authors to make greater efforts to keep the content up-to-date.
In Fig. 7 (a), we observe a rapid growth trend in the development of surveys within the field of “language modeling” since 2023, which is reflected by the increasing number of surveys with positive $IEI$ values. Fig. 7 (b) further supports this finding. Since 2023, the median $RUI$ for language modeling surveys has noticeably exceeded that of surveys in the recommendation systems field. As $RUI$ is positively correlated with the volume of publications within a field, a higher $RUI$ suggests a larger publication volume. While this trend underscores the swift progress in the language modeling domain, it also highlights the growing burden on authors to produce and maintain high-quality surveys in such a rapidly evolving and highly competitive area.
By analyzing the statistical characteristics of numerous review articles’ IEI and RUI, we can uncover developmental patterns across various fields. These patterns provide valuable insights for gauging the momentum of progress within each field and advancing exploration in emerging areas.
5 Indicators-Driven Approach for Screening Reviews
As the number of published reviews continues to grow, it has become common for multiple reviews to cover similar areas, attempt to answer analogous research questions, and often arrive at comparable conclusions. This overlap creates information redundancy and, to some extent, reduces research efficiency. To enhance research efficiency, scholars often rely on academic search engines such as Google Scholar, Semantic Scholar, and Web of Science. They tend to prioritize high-ranking results, typically sorted by citation counts, to identify a select number of highly relevant review articles for detailed examination and analysis. However, these results are sometimes perceived as potentially biased or manipulated [47, 76]. Additionally, an over-reliance on citation counts as a metric can intensify the Matthew effect, which may ultimately hinder sustainable academic development within the field. A feasible alternative is to draw on a set of evaluation indicators rather than relying exclusively on citation counts to inform the selection process.
5.1 Appraising Reviews across Various Topics
A fundamental need may lie in the assessment of reviews across various fields, as most reviews are primarily aimed at readers who are relatively new to a field and may lack an in-depth understanding of the field. For instance, a researcher working in a popular field like object detection may start exploring a less familiar subfield such as optical character recognition (a subfield of computer vision). When encountering a review with only 100 citations, the researcher might dismiss its importance, unaware that in a niche area, 100 citations can actually indicate significant influence. This could lead to an oversight of critical insights in the subfield, simply due to differing citation norms between areas. The proposed $TNCSI$ offers an effective solution by automatically identifying the field of a given article and dynamically estimating its normalized impact based on the field and its current citation count. Illustrative examples are presented in Tab. IV.
TABLE IV: Example Use Cases of $TNCSI$: A similar number of citations (Cites) in various topics may lead to notably different $TNCSI$ values.
Title Cites Topic $TNCSI$
Involvement of Machine Learning for Breast Cancer Image Classification: A Survey [89] 104 Breast Cancer Image Classification 0.79
Class-Incremental Learning: A Survey [152] 99 Class-Incremental Learning 0.61
A Systematic Survey of Prompt Engineering on Vision-Language Foundation Models [35] 100 Prompt Engineering 0.94
End-to-End Speech Recognition: A Survey [96] 100 End-to-End Speech Recognition 0.59
5.2 Tracking Citation Momentum with IEI
Although $TNCSI$ enables cross-domain comparisons, it fails to track changes in citation trends. In such cases, the $IEI$ can offer additional insight. A higher $IEI$ suggests that the paper’s rate of gaining new citations has been increasing over time, reflecting a growing level of recognition within the academic community. Conversely, an $IEI$ below zero implies that the paper’s citation rate has been declining recently. Table V provides examples of $IEI$ applications. In some cases, selecting a paper with a higher $IEI$ may be preferable to choosing one with a higher $TNCSI$, as it reflects a more recent consensus among researchers.
TABLE V: Example Use Cases of $IEI$: A positive $IEI$ indicates that the paper is gaining increasing attention.
Title Cites $TNCSI$ $IEI$
Knowledge Distillation and Student-Teacher Learning for Visual Intelligence: A Review and New Outlooks [121] 565 0.99 -1.41
A Survey on Knowledge Distillation of Large Language Models [132] 40 0.30 1.10
A Review of Practical AI for Remote Sensing in Earth Sciences [49] 29 0.11 -0.16
Change Detection Methods for Remote Sensing in the Last Decade: A Comprehensive Review [23] 23 0.23 0.40
5.3 The Role of Reference Quality in Selecting Reviews
In Sec. 4.3, we briefly discuss the importance of references, noting that the quality of references shows a moderately strong and significant positive correlation with the scholarly impact of a review. Thus, $RQM$ functions as an intuitive metric, allowing researchers to efficiently gauge the average quality of references within a review. Based on the statistical patterns shown in Fig. 6, reviews with an $RQM$ greater than 0.8, an $ARQ$ within the range of 0.7 to 0.9, and an $S_{MP}$ not exceeding 10 are more likely to achieve greater academic impact. Examples of $RQM$ can be found in Tab. VI.
TABLE VI: Example Use Cases of $RQM$: A higher $RQM$ reflects higher scholarly impact ($ARQ$) and a younger age ($S_{mp}$) of the references.
Title $ARQ$ $S_{mp}$ $RQM$
A Survey on Video Diffusion Models [131] 0.72 2 0.94
Video Diffusion Models: A Survey [81] 0.83 3 0.95
A Survey on Multimodal Large Language Models [137] 0.83 1 0.99
Multimodal Large Language Models: A Survey [128] 0.69 5 0.65
A Survey on Benchmarks of Multimodal Large Language Models [57] 0.52 2 0.85
A Survey on Evaluation of Multimodal Large Language Models [43] 0.19 2 0.63
5.4 Assessing the Need for Updating Reviews
Literature reviews inevitably age, yet the rate of aging varies significantly across fields (as shown in Fig. 7). The $RUI$ allows us to gauge the degree of aging of a review. Typically, fields with higher publication rates undergo faster advancements (indicated by a higher $RUI$), meaning that reviews in these areas tend to age more quickly. Conversely, in more established fields with a lower $RUI$, the aging process is correspondingly slower. Compared to relying solely on publication time to assess recency, $RUI$ provides a more accurate estimation. Examples illustrating the exploitation of the $RUI$ are provided in Tab. VII.
TABLE VII: Example Use Cases of $RUI$: $RUI$ quantifies the urgency of updating reviews across different fields.
Title Year Topic $RUI$
A Review of Object Detection Based on Deep Learning [130] 2020 Object Detection 24.10
A Comprehensive Review of Object Detection with Deep Learning [51] 2023 Object Detection 6.72
Time-Series Forecasting with Deep Learning: A Survey [61] 2020 Time-Series Forecasting 29.09
Long Sequence Time-Series Forecasting with Deep Learning: A Survey [22] 2023 Time-Series Forecasting 6.44
5.5 Integrated Use of the Proposed Indicators
In general, researchers are encouraged to apply a set of metrics, rather than relying solely on citation counts to select review articles. In practice, reviews with a $TNCSI$ close to 1 are often prioritized, as this suggests that the review is cited more frequently than most other papers in the same field, indicating widespread esteem within the field. However, supplementary indicators can provide further insights: the $IEI$ captures shifts in citation trends, the $RQM$ evaluates the quality of references, and the $RUI$ reflects how recently the review may require updating. Leveraging a diverse set of indicators allows for the efficient filtering of numerous similar reviews while mitigating the limitations inherent in single-metric citation reliance. More examples of integrated indicators usage are provided in Appendix A.5.
Although the proposed metrics provide an intuitive numerical indication, it is important to emphasize that no existing metric, including the proposed ones, can truly capture the intrinsic value of a review (particularly for systematic review). The proposed impact or quality indicators should be regarded as supplementary instruments rather than definitive measures of quality. It is our contention that the true value of a review resides in its capacity to address the specific research questions and meet the needs of its intended audience. Further explanations regarding the fair use of metrics and how to avoid their misuse are detailed in Appendix A.6.
6 AI-generated Literature Reviews
6.1 Overview of AI-generated Literature Reviews
Traditionally, literature reviews have been manually conducted by researchers who analyze and synthesize scholarly sources to provide an overview of the current state of knowledge in a specific field. Recently, with the advancement of AI technologies, there’s been a growing interest in leveraging artificial intelligence techniques, especially LLMs, to automate or assist in the generation of the literature review [26, 114, 124, 56, 42]. Typically, users are simply required to indicate their area of research interest, and the system will then automatically generate a literature review.
The automated creation of a literature review is a multidisciplinary endeavor that integrates knowledge from various fields. It relies not only on artificial intelligence technologies but also requires the merging of knowledge from other fields such as data science, bibliometrics, database engineering, etc. These components are instrumental in extracting, storing, and synthesizing relevant information from vast amounts of literature. Early attempts in AI-generated literature reviews involve training language models, e.g. LLaMA [116], on a large corpus of academic papers, research articles, and other scholarly content. These models can then be used to generate coherent and contextually relevant text based on prompts or queries related to a specific research topic. However, given that a fully trained or fine-tuned language model has no access to the latest scholarly advances, these models would not generate a literature review that includes the latest scholarly materials. Thus, most existing advanced AI-generated literature review systems contain three main steps: knowledge retrieval, synthesis, and report.
The generation of literature review begins by gathering knowledge from various sources. There is no doubt that a high-quality knowledge retrieval procedure provides a richer context for the LLMs, resulting in the production of more accurate and decent responses. The system usually employs several predefined criteria such as keywords, publication date ranges, and citation numbers to locate, filter, and retrieve relevant resources to be synthesized. These resources are not limited to academic publications, but may also include blogs, official tutorials, and GitHub repositories. Once the relevant knowledge is retrieved, the AI system aims to synthesize the gathered information to create a coherent and comprehensive literature review. This step involves analyzing the retrieved content, identifying relationships between different sources, and organizing the information into meaningful sections. The system may employ LLMs and elaborate prompts to generate the text for each section or the entire survey at once. Finally, the system seamlessly transitions into the report generation stage, where a formatted review is crafted in a cohesive and well-structured style. This stage involves converting the synthesized content into a final report which is ready for presentation. One of the simplest and most common forms of presentation is plain text. For certain systems, the generated text is arranged in various sections based on the predefined setting. Such a multi-step procedure not only streamlines the literature review generation workflow but also ensures the production of acceptable content.
Though researchers are expecting to save significant time and effort in conducting comprehensive surveys by embracing AI-generated literature reviews, numerous concerns about ethical issues and fabrications in AI-generated academic content have been raised. A critical appraisal by Zybaczynska [153] highlights that current AI systems typically fall short in providing substantial, accurate information and critical discernment. Elali et al. [29] critically examines the profound challenges brought by the fabrication and falsification of AI-generated research, underscoring its significant impact on the scientific community. Furthermore, numerous academic publishers are cautious about content created by AI. For instance, journals such as “Nature” and “Science” prohibit listing ChatGPT or other automated tools as authors on papers published in their issues. Nevertheless, we argue that the automated generation of literature reviews not intended for publication remains meaningful. It may assist researchers in staying rapidly abreast of the latest advancements in thriving fields.
6.2 Current State of AI-generated Literature Review
In this subsection, our investigation delves into various efforts that have been made in the field of employing AI techniques to generate literature reviews.
Most of the existing systems are only capable of generating literature reviews in plain text form and with a relatively fixed layout. A popular Github repository ChatPaper [50] retrieves papers based on a user-defined keyword with the use of the Google Scholar crawler. After a cosine similarity-based filtering of the retrieved papers, the selected papers are processed sequentially using ChatGPT. The output of Chatpaper is brief and concise, which usually contains a narrative description of the selected papers. Paper Digest [94] is an AI-based online platform that provides the service of automated review generation. Users can generate a literature review by setting specific keywords and further refining the cited source through constraints on publication dates. Similar to ChatPaper, the output of Paper Digest is also a narrative plain text description. Not all systems for generating literature reviews are fully automated. For example, Jenni AI [2] requires user interaction through “AI Command” during the review creation process, enabling the generation of highly customized literature reviews. Additionally, there are numerous online platforms [105, 14] and plugins available in the GPT Store that offer automated literature review generation services. Given the constraints on manuscript length, detailed introductions to these systems will not be provided.
Despite previous discussions indicating that certain publishers refuse to accept non-human authors, scholars have still successfully published AI-generated literature reviews in some of academic journals. Aydın et al. investigate how well can LLMs perform in the generation of the literature review. They employed well-known LLMs including ChatGPT and Google Bard to generate reviews on digital twin in healthcare [7] and meta-universes [6]. In their attempts, the Google Bard (or the ChatGPT) was first adopted to paraphrase the abstracts of papers within the last 3 years. Then, they designed several question prompts (such as “What is digital twin”) to query the LLMs and rearranged the response in a formatted style.
Although significant progress has been made in AI-generated literature reviews, there is still substantial room for improvement. Achieving automated literature reviews that match the quality of human-authored ones remains an active area of research, highlighting the need for further exploration and development. For instance, most current systems for generating literature reviews cannot extract information from tables or figures, resulting in the exclusion of these elements from the AI-generated reviews. Further details on the distinctions between AI and human-authored reviews are provided in the next subsection.
6.3 Human-authored vs. AI-generated
We first provide a descriptive analysis aimed at elucidating the distinctive approaches employed by human researchers and their AI counterparts. To prevent potential information leakage and ensure a fair analysis due to the possibility of ChatGPT having been exposed to relevant reviews during training, we selected a newly emerging field, i.e. visual prompt engineering, as the focus of our investigation. Table VIII reveals that nearly all automatic summarization systems fail to incorporate visual elements, which are highly valuable for readers. Moreover, most automated review systems lack explicit appraisal criteria, making it challenging to ensure the quality of the references included.
Review Refs Auto-Level Vis. Elems. SALSA Analysis [33]
Human [119] 160 Manullay A typically narrative, method clustering-based review aims to offer valuable insights in visual prompt learning. The selection criteria for references are not specified, but it’s clear that each reference has been thoroughly appraised.
Jenni [2] - Semi-automated $\times$ Automated searching and appraising references are not supported. The generation process relies on user interaction, and only a narrative description is provided.
AutoSurvey [124] 143 Automated $\times$ Narrative review without any visual elements. Although no formal appraisal framework is utilized, the approach incorporates semantic verification and multi-LLM evaluation. Generated reviews include narrative descriptions of key contributions, research trends, and gaps.
ChatPaper [50] 15 Automated $\times$ Retrieving relevant literature from arXiv based on multiple LLM-generated keywords without appraisal step. Each section contains plain descriptions of the related reference.
PaperDigest [94] 20 Automated $\times$ Searching papers with the user-specified keyword. It seems to appraise the quality of references with private criteria. The generated content is more like a concise summary.
askyourpdf [14] 9 Automated $\times$ No official explanations are found for how to retrieve references and appraise the quality of the literature. The generated review includes a brief analysis and description of the related concept, current state of development, and gaps.
TABLE VIII: Comparisons between human-authored and AI-generated literature reviews. “Vis. Elems.” denotes the visual elements.
In Fig. 8, we further present a set of scatter plots depicting the references’ quality of (a) reviews published in several academic journals and (b) reviews generated by various AI systems. To allow sufficient time for the references selected by the AI-generated review system to accumulate academic impact, a two-month interval is set between obtaining the reference list and calculating their $TNCSI$. The horizontal axis represents the reference age, and the vertical axis corresponds to the proposed $TNCSI$. The color and size of the scatter points indicate the $IEI$ of the references. A positive $IEI$ value signifies a gradually increasing citation trend, resulting in warm-colored scatter points. The larger the size of the scatter point, the greater the value of $IEI$. Conversely, when the $IEI$ value is negative, the scatter points are cool-colored, and the size decreases as the value decreases. A closer examination of Fig. 8 reveals a distinct distribution among human-authored and AI-generated depicted reviews in their respective scatter plots. Human-authored reviews exhibit clustered dots in the upper left corner, indicating a trend toward referencing more recent and influential sources (i.e., citing newly published works with limited influence at the time of writing). Conversely, AI systems predominantly cite established, high-impact papers, creating a relative sparsity in the upper left corner. We argue that this discrepancy stems from seasoned researchers’ ability to assess a paper’s value through a diverse, fine-grained, and content-based manner, rather than relying solely on citation counts. This allows them to render more precise judgments about the potential significance of a paper. In contrast, most existing AI-based review systems depend exclusively on citation counts as their primary filtering metric, limiting their capacity to identify high-quality recent publications as references. We also note the presence of a potential boundary line for the $IEI$ metric in Fig. 8, panel (b), which is likely a result of the quality control mechanisms in certain automated review systems.
Figure 8: Visualization of references quality of human-authored and AI-generated literature reviews: Panel (a) illustrates the quality-age distribution of references in human-written reviews, while panel (b) depicts the quality-age distribution of references in AI-generated reviews. It can be observed that the TNCSI of references less than one year old in AI-generated reviews is significantly lower than that in human-authored reviews. This suggests that current AI-generated review systems struggle to appraise the academic value based on the article content.
In summary, the utility of AI-generated reviews remains limited for several reasons. Firstly, they often lack adequate visual elements and adversely affect the overall readability and user experience. Secondly, current systems face challenges in accurately assessing the quality of newly published articles, which undermines the key advantage of AI-generated reviews—timeliness. At this stage, users are encouraged to review the selection criteria of AI-generated review systems before using them and carefully assess the newly published references included. Additionally, we would like to emphasize that researchers should fully leverage another significant advantage of AI-generated reviews: high customizability. Researchers could provide clear and specific research questions when using AI review systems to ensure the generated reviews meet their specific needs.
7 Challenges and the Future of the Literature Review
Challenges and future opportunities for both human-authored and AI-generated reviews are discussed in this section.
7.1 Challenges
Information Overload. While our proposed indicators help alleviate information overload to some extent, several challenges remain unresolved. First, As scientific research continues to expand rapidly, the volume of literature review is expected to grow significantly. Collecting and screening a vast number of publications in these emerging fields will likely lead to incomplete searches. Both human authors and AI systems are required to address the challenge of identifying the most relevant references from the extensive body of published literature. In addition to incomplete searches, the research community should consider exploring ways to avoid redundant efforts. For example, the first two review papers [143, 144] on the segment anything model [53], uploaded to arXiv just two days apart, shared at least 70 overlapping references. Avoiding such duplication not only helps reduce the waste of resources (despite both papers offering unique insights) but also minimizes information redundancy.
Lag and Obsolescence.
Scientific knowledge is constantly evolving, yet the process of compiling reviews often struggles to keep up with the latest research advancements—particularly when created by human authors and subject to peer review. Comprehensive reviews are often delayed relative to the emergence of new research areas. For instance, when Mamba [34] started gaining traction, no in-depth reviews were available during its initial months. This delay reflects the time-intensive nature of gathering, evaluating, and synthesizing an ever-expanding body of literature, along with the requirements of the peer-review process.
7.2 Future
AI Empowerment Literature Review. While debates persist regarding the ethical implications and appropriateness of AI-generated content for formal academic publication, such AI-driven reviews have gained acceptance for informal purposes, e.g., rapid ad-hoc reviews [101]. Considering the substantial time and effort required for comprehensive literature reviews, leveraging AI to automate review generation presents a promising opportunity that deserves further exploration. The future development of AI-generated reviews could focus on achieving more precise content-based evaluations of article quality and integrating visual elements to enhance their utility and readability.
Long-term Support Literature Review. The concept of long-term support (LTS) for literature reviews offers a practical solution for maintaining relevance and timeliness in the fast-paced environment of academic research. Unlike traditional reviews, which are generally published once and remain static, an LTS literature review would involve regular updates to include new research findings and advancements in the field. This approach is especially valuable when authors have invested considerable effort in developing structured frameworks or taxonomies that serve as foundational tools in their domain. Such frameworks often require the author’s expertise to adapt effectively to new findings while preserving the original rigor and intent. By maintaining their reviews, authors may ensure that these foundational classifications evolve responsibly, allowing the field to benefit from both continuity and the latest advancements.
8 Conclusion
In conclusion, this paper conducts a tertiary study on the extensive body of literature reviews in the PAMI field to address three core research questions: (1) What are the structural and statistical features commonly observed in PAMI literature reviews? (2) What strategies can researchers adopt to efficiently navigate the growing number of literature reviews in PAMI? and (3) How can the differences between AI-generated and human-authored reviews be understood to ensure their effective use while mitigating potential risks? To tackle these questions, we perform statistical analysis using the RiPAMI database, and introduce four article-level bibliometric indicators to evaluate review papers from various perspectives.
For the first research question, we conduct an analysis and synthesis of reviews from various disciplines, identifying the dominant structural patterns and shared features specific to PAMI reviews. We develop the RiPAMI database, which contains metadata for over 3,000 reviews. Statistical analysis of this database allows us to pinpoint key statistical features related to word count, reference count, and other metadata elements in PAMI reviews. Furthermore, we utilize the LLM for information extraction, demonstrating that no more than 20% of reviews in the PAMI field comply with PRISMA standards, although this percentage gradually increases each year. We also reveal the positive correlation between the reference quality and academic impact, and examine how the proposed indicators offer insights into the field development trends.
For the second research question, we discuss how the proposed four bibliometric indicators assist researchers in screening review papers. The proposed indicators offer an innovative alternative to traditional bibliometric analysis, enabling a cost-efficient, field-normalized, and real-time assessment of literature reviews. These bibliometric indicators not only furnish clear numerical hints for human researchers but could also offer substantial assistance in the appraisal processes of AI-generated review systems.
For the last research question, this study underscores a significant distinction between human-authored and AI-generated literature reviews: AI-generated reviews typically face challenges in evaluating the potential value of newly published articles. The insights gained from this comparison provide a guide for the effective use of automated review systems and point the way forward for the future development of these systems.
To further support and expand upon this work, we release the code framework that encompasses functionalities including metadata retrieval, indicator calculation, and data analysis, etc. While it is designed for analysis of literature reviews in PAMI, the framework is believed to be universally adaptable, catering to the expansive needs of researchers from disparate fields.
Appendix A Details of the Proposed Indicators.
In this section, we provide further insights into the proposed indicators. As presented in Tab. IX, various previous efforts have aimed to tackle the challenge of comparing metrics across fields. For instance, Paper [97] analyzes two well-known field-independent citation metrics at the article level: Field-Weighted Citation Impact (FWCI) and Relative Citation Ratio (RCR). This study suggests that both metrics perform similarly in normalizing citations across research domains. However, these metrics require predefined field categorizations (such as the Scopus All Science Journal Classification) and may not adequately assess papers in emerging subfields. The Field Normalized Citation Success Index (FNCSI), introduced in Refs. [109, 115], is defined as the likelihood that a paper published in Journal A is cited more frequently than a randomly selected paper from Journal B. Although FNCSI offers robustness, it should be noted that it depends on predefined topic keywords and is specifically suited for journal-level assessments.
As the Leiden Manifesto [41] states: metrics should “Account for variation by field in publication and citation practices”. We develop the concept of impact indicators, which is a measure used to gauge the impact of a certain paper in its field. The reason for using the term “indicator” rather than “metric” is we believe that the impact of a certain paper cannot be fully and accurately measured. While metrics like citation counts, h-index, or journal impact factors could indicate a paper’s influence within the academic community, they fail to capture all aspects of its impact. For instance, a research paper might lead to significant advancements in theory, methods, or understanding in its field, none of which would necessarily be reflected in academic metrics. Similarly, a paper might contribute novel concepts or techniques that become influential over time but are not initially evident in citation counts. Hence, the term “indicator” is favored, as it suggests a signal without asserting to encompass the entirety of the academic paper’s contribution.
Metric Assessing Level Normalized Pre-defined Keywords Free
Citation Counts A/J $\times$
Impact Factor J $\times$ $\times$
FNCSI [115] J Field and Value $\times$
CiteScore [113] J $\times$ $\times$
SNIP [87] J Field $\times$
FWCI [31] A Field $\times$
RCR [46] A Field $\times$
TNCSI(Ours) A/J Field and Value
TABLE IX: Metrics for Evaluating Scholar Impact of Papers: “A” and “J” stand for article-level and journal-level. Field normalized signifies that the metric can be utilized across fields. Value normalized indicates that the range of the value is between 0 and 1.
A.1 TNCSI
$TNCSI$ is a metric used to assess the academic impact of articles across various disciplines. The fundamental distinction of $TNCSI$ from other metrics lies in its independence from the predefined keyword. The requirement for a predefined keyword is a primary reason that most bibliometric indicators face challenges in facilitating cross-disciplinary comparisons. To address such a limitation, we propose adopting ChatGPT [91] (gpt-3.5-turbo-0125) to generate the topic keyword and retrieve related papers according to the keyword. As illustrated in Fig. 9, we ask ChatGPT to identify the most representative topic keyword with the paper’s title and abstract. ChatGPT is one of the most advanced and influential LLMs in the field of natural language processing [149]. Equipped with state-of-the-art language understanding capabilities, ChatGPT has revolutionized the way we interact with AI-powered conversational systems by simply setting “system”, “user”, and “assistant” roles. The “system” role sets the conversation’s behavior and initial context. It provides instructions to guide the assistant’s responses. The “user” role represents the individual interacting with ChatGPT, who inputs messages to the assistant. The “assistant” role is the ChatGPT model itself which would respond based on the provided instructions and user input.
Figure 9: Conceptual illustration of topic keyword generation process. Few-shot prompting may enhance the response quality of the large language model.
We count the citations of papers with the help of Semantic Scholar API. For each paper, we retrieved up to $k=1000$ relevant sources using the API. Based on the selected $k$ papers and the corresponding citation counts $p_{c}$ for each paper $p$, we can calculate the discrete citation frequency distribution of the $k$ papers in a certain topic. We may further consider the distribution as a probability mass function:
$P(X=x)=\begin{array}[]{ll}\frac{{\text{Citation}_{x}}}{k},\end{array}$ (5)
where $Citation_{x}$ represents the number of papers with $x$ citations. The $k$ papers related to the topic and their meta-data could be retrieved using a pre-defined topic keyword through online scholar search engines or API, such as Semantic Scholar, CrossRef, or Google Scholar, as mentioned in Sec. 2.3.
Considering that the number of papers with $x$ citations generally follows an exponential decay, we utilize the maximum likelihood estimation method to fit $P(X=x)$ and obtain the probability density function (PDF) of a continuous exponential decay distribution:
$f(x)=\lambda e^{-\lambda x},x\geq 0.$ (6)
Next, we will calculate $TNCSI$ by performing a definite integral on $f(x)$.
[…]
where $f(x)$ represents the probability density at the value $x$, and $\lambda$ is the results obtained from the maximum likelihood estimation, representing the scale parameter controlling the scaling. Finally, the definite integral of $f(x)$ over the interval $[0,citeNum]$ gives us the desired $TNCSI$:
In most scenarios, the generated topic word is sufficient to meet expectations, which can be further used as the keyword to retrieve papers from online scholar search engines. Optionally, one can set “System”, “User”, and “Assistant” roles before the final query to improve response quality and create more tailored interactions with the ChatGPT. In other words, a few-shot user-assistant pair prompts the ChatGPT with context on topic granularity. For example, a paper about the classification of irises by an improved CNN may have different perspectives. Some researchers focus more on the algorithm of the improved CNN, while others may be interested in classifying irises. Such ambiguity would likewise make it difficult for ChatGPT to identify the most representative topic keyword as expected. However, this could be addressed by providing the context which consists of (1) the identical prompt template with the replaced title and abstract as user input, and (2) the expected topic keyword as assistant output. By default, we adopt a well-known paper [27] as an example to guide models to generate the topic keyword for all papers.
Similar to other LLMs, ChatGPT adapts its responses based on user-provided natural language prompts. However, natural language prompts can be ambiguous, and different prompts may lead the model to respond with varying quality. Thus, we follow the practices of LLM prompt engineering and carefully design the prompt to optimize the desired output. To determine the optimal prompt, we construct a dataset by manually annotating the topic keywords of 201 papers from various domains published later than the ChatGPT being trained and then compare the performance of multiple prompts on this dataset. The normalized edit distance [141] is adopted to measure the similarity between the GPT-generated keyword and our annotated keyword, where a lower value indicates a higher quality of the prompt. As can be seen from Tab. X, some of the designed prompts achieve decent NED scores for papers in various domains.
NO. User Prompt Content Few-shot NED↓
1 Please analyze the title and abstract provided below and identify the main topic or central theme of the review paper. Focus on key term and the overall subject matter to determine the primary area of research or discussion.The output should be formatted as following: xxx $\times$ 0.75
2 Given title and abstract, please provide the searching key phrase for me so that I can use it as keyword to search highly related papers from Google Scholar or Semantic Scholar. Please avoid responding with overly general keyword such as deep learning, taxonomy, or surveys, etc. Answer with the words only in the following format: xxx $\times$ 0.40
3 Identifying the topic of the paper based on the given title and abstract. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that is unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.36
4 Identifying the topic of the paper based on the given title and abstract. So that I can use it as keyword to search highly related papers from Semantic Scholar. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that is unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.32
5 Identifying the topic of the paper based on the given title and abstract. I’m going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that are unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\times$ 0.29
6 Identifying the topic of the paper based on the given title and abstract. I’m going to write a review of the same topic and I will directly use it as keyword to retrieve enough related reference papers in the same topic from scholar search engine. Avoid using broad or overly general term like ’deep learning’, ’taxonomy’, or ’surveys’. Instead, focus on keyword that are unique and directly pertinent to the paper’s subject. Answer with the word only in the following format: xxx $\checkmark$ 0.28
TABLE X: Effectiveness of prompt engineering: The few-shot approach slightly improves the performance of topic keyword extraction.
To provide a clear understanding of the proposed quality indicators, We render the graphical representations of $TNCSI$ in Fig. 10. As can be seen in Fig. 10, $TNCSI$ equals the area under the probability density function curve, which is fitted with the use of maximum likelihood estimation.
Figure 10: Visualization of the proposed $TNCSI$. The $TNCSI$ equals the area under the fitted probability density function curve.
A.2 IEI
The $IEI$ is an indicator used to measure changes in citation trends over a given period. A direct method for evaluating changes in citation trends is to apply polynomial fitting to the scatter data and calculate the sum of the derivatives at each point. However, in practice, polynomial fitting methods tend to be sensitive to outliers when applied to data lacking discernible distribution patterns, potentially compromising their robustness significantly. Consequently, the numerical values obtained may not accurately reflect the underlying citation trend. In contrast to these series analysis-based methods, we propose a morphological theory-grounded Impact Evolution Index ($IEI$), which converts the citation trend into a clear and interpretable numerical value.
To calculate the $IEI$, we first need to obtain the distribution of citation counts over time since publication. Once the citation data is retrieved, we may create a sequence $Seq_{\text{citation}}$ about the number of citations. The $i\in\{0,1,2...,l\}$ item in the sequence $Seq_{\text{citation}}[i]$ represents the number of received citations in the $i_{\text{th}}$ month after the publication, where $l$ represents the number of months allocated for trend observation. Then, a sequence $Seq_{\text{time}}$ of the same length as $Seq_{\text{citation}}$ is generated by enumerating from 0 to $l-1$, Typically, the minimum recommended value for $l$ is 6 or higher. This ensures that the data used for the analysis is adequately representative and the results are reliable. We match the items at the same positions in $Seq_{\text{time}}$ and $Seq_{\text{citation}}$ to determine a set of discrete coordinates $\{(Seq_{\text{time}}[i],Seq_{\text{citation}}[i])\}$, which serve as the control points for shaping the Bézier curve. A Bézier curve is a mathematical representation of smooth curves commonly used in computer graphics, image editing, and design software. The curve starts at the first control point and ends at the last control point, while the intermediate control points influence the curvature and direction of the curve. The number of control points determines the degree $n=l-1$ of the curve.
$C(t)=\sum_{i=0}^{n}B_{i,n}(t)P_{i},$ (8)
$B_{i,n}(t)=\binom{n}{i}(1-t)^{n-i}t^{i},t\in[0,1],$ (9)
where $B_{i,n}(t)$ represents the coefficient of the Bézier curve at a given parameter value $t$, which determines the position along the curve ($t=0$ means the start and $t=1$ means the end). $\binom{n}{i}$ is the binomial coefficient, also known as “n choose i”. It represents the number of ways to choose $i$ elements from a set of $n$ elements. $P_{i}$ stands for the $i_{\text{th}}$ control point of the curve.
Given the continuity of the Bézier curve, we can compute its derivative as follows:
$C^{\prime}(t)=n\cdot\sum_{i=0}^{n-1}B_{i,n-1}(t)\cdot\left(P_{i+1}-P_{i}\right).$ (10)
The tangent vector $C^{\prime}_{a}$ at the $a$-th point on the Bézier curve is further given by Eq. (11).
$C^{\prime}_{a}=$ $n\cdot\sum_{i=0}^{n-1}B_{i,n-1}(\frac{a}{n})\cdot\left(P_{i+1}-P_{i}\right),$ (11)
$=$ $(x_{a},y_{a}),\ a=0,1,\cdots,n,$
where $x_{a}$ and $y_{a}$ are components of the vector, representing its magnitude along the x and y axes respectively.
Finally, the $IEI_{L_{l}}$ can be obtained by averaging the slope of $l=n+1$ distinct points on the curve, as already shown in Eq. (2). Moreover, different months may contribute differently to the $IEI$. For instance, if we desire closer months to have a greater impact, we can achieve this by adjusting the weighting coefficients, $w_{a}$, of the slope at different points to calculate their weighted averages (See in Eq. (12)). In addition, the instantaneous trend could be regarded as the slope of the last month in the sequence. It can be obtained by setting $w_{n}=1$ and the other weighting coefficients to 0. We denote the $IEI$ focused on the last month among the latest $l$ months (excluding the current month) as $IEI_{I_{l}}$, as depicted in Eq. (13).
$IEI_{W_{l}}=\sum_{a=0}^{n}\frac{w_{a}(y_{a}/x_{a})}{n+1},$ (12)
$IEI_{I_{l}}=C^{\prime}(1)=n\cdot(P_{n}-P_{n-1}).$ (13)
Note that the value of $l$ can be configured flexibly to meet actual demands. In general, the longer the period being analyzed, the more stable the citation trend becomes. We usually prefer to analyze the most recent 6 months of citations when constructing a Bézier curve of degree 5 and calculate $IEI_{L_{6}}$, $IEI_{W_{6}}$, and $IEI_{I_{6}}$ based on the curve.
The illustration of the $IEI$ is shown in Fig. 11. In Fig. 11, the horizontal axis labeled 0 to 5 inversely denotes the months prior to the current month, with 0 representing 6 months ago and 5 denoting the previous month. The corresponding $IEI$ is approximately $-0.76$, which indicates a slightly decreasing citation trend of the investigated review over the last six months.
Figure 11: Visualization of the proposed $IEI$. $IEI$ is the average of the derivatives of each control point on the Bézier curve.
A.3 RQM
The quality of references in a literature review is a multifaceted concept involving various aspects, such as credibility, relevance, breadth, and depth. Quantitatively assessing these elements is challenging. To begin with, precisely defining the relevance of scientific literature is difficult; both co-citation analysis [111] and similarity in paper embeddings have inherent conceptual limitations. For instance, while the concept of breadth in a literature review can theoretically be measured as the ratio of cited references to the total number of relevant references, accurately calculating this ratio is challenging. Identifying all relevant references through keyword searches or citation networks remains elusive.
Given the challenges in quantifying reference quality through the direct factors mentioned, we instead consider an indirect indication of reference quality using $TNCSI$. Here, $TNCSI$ serves as a quality indicator based on collective user assessments, effectively capturing a statistical summary of numerous researchers’ evaluations of these direct factors for a given paper. Timeliness also matters. An up-to-date literature review ensures that the latest advancements, developments, and perspectives within a field are incorporated. This temporal relevance enhances the accuracy and effectiveness of research outcomes by reflecting the current state of knowledge. By emphasizing the currency of references, we can assess how thoroughly the literature review integrates the most recent research and developments in the field.
To account for both the quality and timeliness of cited references, we propose modifying the Gompertz function to model the reference quality (see in Eq. (3)). The Gompertz function is characterized by its sigmoidal, which indicates a slow growth rate at the start and end of a time period, with more rapid growth in the middle phase. This pattern is often observed in natural phenomena, such as species dynamics [16], tumor growth [118], etc.
The value of $RQM$ is determined by the average reference quality (ARQ), the shift parameter $\beta$, and the median age of references $S_{mp}$. The calculation procedures of $ARQ$ are as follows: The first step is the extraction of the cited reference list. For most publications, their reference lists could be provided by Semantic Scholar API. For a small number of reviews, the reference list provided by Semantic Scholar may contain errors. In this case, although there are powerful computer vision-based algorithms [13, 25] available for extracting the reference list within PDFs, our requirements are relatively simple and can be effectively met by relying on the heuristic algorithms or ChatGPT. More specifically, for literature review with a relatively fixed citation format, we can use the PDFMiner [95] to read text from PDF files and use heuristic rules to match citations. Alternatively, the text can be analyzed using ChatGPT to extract in-text citations. The second step is similar to the approach presented in Sec. 2.4.1, where the ChatGPT and a well-designed prompt (as presented in Fig. 9) are utilized to obtain the topic keyword of the review. Next, we calculate the $TNCSI$ for each reference in the list. To conserve computational resources, we avoid using the ChatGPT to generate keywords for each reference. Instead, the $TNCSI$ of all cited literature is calculated using a sharing topic keyword. Finally, the coverage can be further calculated in Eq. (14):
$ARQ=\frac{\sum_{i=1}^{N_{R}}TNCSI(Ref_{i})}{N_{R}},$ (14)
where $TNCSI(\cdot)$ refers to the $TNCSI$ value of the $i_{th}$ cited reference, and $N_{R}$ stands for the number of the reference. In certain instances, it has been noted that calculating $TNCSI_{s}$ for each cited literature is also reasonable. However, this paper primarily emphasizes the current impact of the cited references, hence the utilization of $TNCSI$ in this context.
The shift parameter $\beta$ can be set empirically or obtained statistically. For statistical estimation, we first examine the distribution of $S_{mp}$ and $ARQ$ of the RiPAMI database. The results indicate that (1) the $S_{mp}$ of over 50% of the reviews falls within the [5,10] interval, where we denote the lower boundary as $l_{s}$ and the upper boundary as $r_{s}$; (2) the average of $ARQ$ is approximately 0.6, denoted as $\overline{ARQ}$. Then, the problem of asserting for $\beta$ is reconceptualized as an optimization problem. As shown in Eq. (15), the objective here is to identify the value of $\beta$ that maximizes the integral of the absolute value of $RQM^{\prime}(S_{mp};\beta,\overline{ARQ})$ over the range $l_{s}$ to $r_{s}$, subject to the constraints of $\overline{ARQ}=0.6$. Such an approach would endow $RQM$ with a more discriminative nature. It should be noted that different fields may result in distinct values of $\beta$. For this study, the $\beta$ has been set as 5 for all fields.
[…]
The range of $RQM$ extends from 0 to 1, where values closer to 1 signify a higher quality of the referenced literature. As illustrated in Fig. 12, when the $ARQ$ of a paper remains constant, an increase in the variable $S_{mp}$ will lead to a decrease in the $RQM$ value. Conversely, when $S_{mp}$ remains constant, a higher $ARQ$ will elevate the $RQM$ value.
Figure 12: 3D visualization of the proposed $RQM$
A.4 RUI
To evaluate the $RUI$, we may start with the coverage of references before and after publication. This coverage ratio can, to some extent, indicate the extent to which a review requires updating within its field. However, as mentioned earlier, accessing the coverage of a review is difficult. Fortunately, this problem is subtly avoided in calculating the ratio of relevant papers before and after publication. Assuming that the ratio of references containing the topic keyword in the title to all references is $R_{k}$, the total number of relevant articles can be estimated by dividing the number of articles containing those keywords retrieved from a search engine by $R_{k}$. Note that $R_{k}$ generally remains consistent before and after the publication, the Coverage Difference Ratio ($CDR$) can then be calculated in Eq. (16). The theoretical value range of $CDR$ is greater than 0 to positive infinity. When the $CDR$ of a review equals 1, it indicates that the current field has yielded new publications sufficient to constitute half of the literature referenced in the review.
$CDR=\frac{N_{pc}\cdot R_{k}}{R_{k}\cdot N_{mp}}=\frac{N_{pc}}{N_{mp}},$ (16)
where $N_{mp}$ and $N_{pc}$ denote the number of relevant literature from the median publication date of the cited references to the publication date of the review, and from the publication date of the review to the current time, respectively.
In addition, similar to the inevitable process of biological aging, literature reviews also undergo a gradual aging process throughout time. Such passage of time bestows upon literature reviews increasing aging progress, where the degree of aging can be conceptualized as a normalized value of the academic impact already achieved. To further explore the aging of reviews in the field of PAMI, we conducted a statistical analysis of the yearly number of newly received citations for reviews published between 2015-2017 in RiPAMI. In contrast to earlier findings, however, the distribution of received citations of reviews over time follows a t-distribution rather than a log-normal distribution of the regular paper, as previously reported in Ref. [79, 28, 86]. Due to the insufficient duration of published sample data, the observation of citation-time trend curves is incomplete. Therefore, we conducted a three-degree polynomial fitting on the limited 6-year citation trend data and transformed the positive segment of the fitted curve into a PDF. To obtain the corresponding cumulative distribution function (CDF), we employed the cumulative trapezoidal numerical integration method for an approximate estimation. Thus, the Review Aging Degree (RAD) is given by:
[…]
where $M_{pc}$ denotes the duration in months from the publication of the review to the present, $p=-0.003$, $q=0.001$, $r=0.1267$, $s=0.0129$ are the coefficients obtained by polynomial fitting. Please note that the integral symbol used here is for illustrative purposes only. The strict mathematical definition involves the accumulation of discrete trapezoidal areas.
Finally, the $RUI$ could be obtained by weighted summation of $CDR$ and $RAD$, as shown in Eq. 4. A sculptural visualization of the $RUI$’s contours is crystallized in Fig. 13.
Figure 13: 3D visualization of the proposed $RUI$
A.5 More Examples of Indicator Usage
Tab. XII provides more examples of how the four proposed indicators further assist researchers in refining their review selections. Compared to relying solely on titles, citation counts, and publication dates, the proposed metrics provide more information from various perspectives to assist researchers in review selections.
A.6 Ensuring Fair Use and Avoiding Metric Misinterpretation
No metric is perfect. Researchers should recognize the limitations of metrics and avoid their misuse or misinterpretation. All citation-based metrics, in particular, are subject to various biases, such as those influenced by the Matthew effect, which can inadvertently contribute to the persistence and dissemination of inaccuracies within academic literature [112].
For TNCSI. $TNCSI$ is primarily used to assess the cumulative impact of literature reviews and should not be used to evaluate the quality of literature reviews, especially those published in different years.
For IEI. $IEI$ solely represents citation trend changes over a specific past period and should not be extrapolated to predict future citation trends solely.
For RQM. As mentioned earlier, due to practical and conceptual constraints, $RQM$ considers only the average impact and timeliness of references, without accounting for the relevance of the references to the investigated topic. Consequently, there is a potential for manipulating this metric by citing influential references unrelated to the topic to inflate the $RQM$ score.
For RUI. As its name suggests, the $RUI$ is a metric used to assess the extent to which a review requires updating. It is not designed to quantify the degree of obsolescence of the perspectives or conclusions within the review.
Additionally, it is important to note that all the metrics are designed for article-level analysis of reviews and should not be used to infer the development of an entire field based on the metrics of a single article.
A.7 Alignment with the Leiden Manifesto
Principle $TNCSI$ $IEI$ $RQM$ $RUI$
Quantitative evaluation should support qualitative, expert assessment.
Measure performance against the research missions of the institution, group or researcher. $\times$ $\times$ $\times$ $\times$
Protect excellence in locally relevant research. N/A N/A N/A N/A
Keep data collection and analytical processes open, transparent and simple.
Allow those evaluated to verify data and analysis.
Account for variation by field in publication and citation practices.
Base assessment of individual researchers on a qualitative judgement of their portfolio. N/A N/A N/A N/A
Avoid misplaced concreteness and false precision.
Recognize the systemic effects of assessment and indicators.
Scrutinize indicators regularly and update them. N/A N/A N/A N/A
TABLE XI: Alignment of Leiden Manifesto Principles with Proposed Metrics: “✓” indicates alignment with the principle, “$\times$” indicates non-alignment, and “N/A” indicates principles not applicable.
Table XI shows the alignment of our proposed four metrics with the specific guidelines of the Leiden Manifesto. The metrics—$TNCSI$, $IEI$, $RQM$, and $RUI$—demonstrate strong adherence to most principles outlined in the manifesto. Specifically, all four metrics align with principles emphasizing the integration of quantitative evaluation with qualitative expert assessment, transparency in data collection and analysis, verification of data and analysis by those evaluated, accounting for field-specific variations, and avoiding misplaced concreteness and false precision.
The metrics do not align with the principle of measuring performance against the research missions of the institution, group, or researcher, as indicated by the “$\times$” in the corresponding row. We would like to point out that this reflects a deliberate trade-off made to balance engineering complexity. While more fine-grained measurements of performance would indeed enhance the objectivity and fairness of the metrics, they would inevitably increase the difficulty of data acquisition. In the future, with advancements in academic search engines and breakthroughs in natural language understanding, this issue is expected to be effectively addressed.
TABLE XII: Comparison of various reviews with the proposed indicators and metrics.
[…]
Appendix B Information Extraction Techniques
B.1 Word Counts
Word counting differs from character counting, where methods relying solely on regular expressions often yield imprecise results due to the complexity of word segmentation rules. To achieve precise word counting, we utilize the Python NLTK library for word tokenization and exclude all tokens representing non-alphabetic elements. As a result, the presented word count only includes alphabetic words, excluding numbers, contractions, and any special symbols.
B.2 Visual Elements Counts
[…]
Figure 14: LLM for figure and table caption extraction: This LLM RAG-based approach allows for the retrieval of figure and table captions at an acceptable cost.
B.3 Review Feature Extraction
This paper performs a statistical analysis of six features across more than 3,000 review samples, providing insights into factors such as compliance with PRISMA standards and the presence of application sections. Similar to visual element counting, these features are identified using LLM RAG-based information extraction methods.
Our proposed method for extracting review features is depicted in Fig. 15. The process begins with document analysis and recognition. Due to the inherent structure of PDF files, the machine-readable text order often diverges from the natural reading sequence. Additionally, elements such as captions and superscripts can introduce noise, disrupting semantic continuity during direct text extraction. To overcome these challenges, we propose utilizing the Nougat model [13] to process document images and convert them into structured text. This structured text enables efficient extraction of the Table of Contents (TOC) and content from individual sections.
Following the discussion in Sec. 3.1, we adopt a cost-effective approach during RAG by limiting the need for the LLM to process the entire text. In simple terms, specific sections of the article are selectively provided to the LLM to generate targeted responses. Alongside the title and abstract, the extracted introduction section and TOC are provided to guide the LLM in identifying whether the authors propose a new taxonomy or include a dedicated section for inclusion and exclusion criteria (indicating adherence to PRISMA guidelines) in the reviewed literature. To assess whether the authors discuss preliminaries, applications, or future challenges, we propose relying exclusively on TOC information, further minimizing costs. For benchmarking, the LLM is instructed to analyze extracted captions (illustrated in Fig. 14) to determine whether the authors have performed quantitative benchmarking of existing methods. All responses generated by the LLM are formatted into a structured 0/1 schema and stored in the RiPAMI database.
Figure 15: LLM for review features extraction: By utilizing OCR technology [13], we extract structured text and input the introduction section (Intro.), table of contents (TOC), and captions of visual elements into the LLM to derive review features.
Appendix C The Relationship between Review Features and TNCSI
In this study, we dedicate our efforts to analyzing and synthesizing six key types of content features in review articles within the PAMI field. It is important to emphasize that the presence or absence of these features in an article should not be interpreted as a direct measure of its overall quality or potential academic value. As shown in Fig. 16, there is no apparent correlation between the distribution of these six features and $TNCSI$.
Figure 16: Analysis of six review content features and their association with $TNCSI$: Gene expression heatmap illustrating the independence between six review content features (PRISMA, Application, Benchmark, Preliminary, Taxonomy, and Discussion) and $TNCSI$. Darker shades indicate a match with the feature, corresponding to a value of 1.
Appendix D Code Framework: PyBiblion
This study involved extensive coding and engineering practices during the processes of dataset construction and statistical analysis. To support this work, we developed an open-source code library, PyBiblion, designed for bibliometric and statistical analysis. As its name suggests, PyBiblion is a Python-based library designed for bibliometric and statistical analysis, drawing from the Greek root “biblion”, meaning “book” or “document” to emphasize its focus on scholarly literature.
PyBiblion offers several core advantages, one of which is lazy loading technology. Unlike most existing frameworks, PyBiblion implements lazy loading for paper information. This means that no network communication occurs during the initialization of an instance object; instead, data is only loaded when it is explicitly accessed. This mechanism helps to reduce the number of requests, alleviating server-side communication pressure and fostering a more equitable usage environment. The second advantage of PyBiblion lies in its user-friendliness. Extensive engineering optimizations have been implemented to integrate information retrieval and metric computation seamlessly. This design enables users to compute metrics such as the proposed $TNCSI$ and $RQM$ with a single line of code, offering a highly efficient and intuitive experience. In addition, PyBiblion integrates numerous practical features, including database support, multithreading execution, visualization tools, and statistical analysis capabilities, further enhancing its functionality and versatility.